OpenStudy (anonymous):

A lawn mower manufacturer incurs a total of $34,816 in overhead costs and $388 per lawn mower in production costs. How many lawn mowers were manufactured if the average cost of production is $660? The number of lawn mowers manufactured was ___?

OpenStudy (mikeymaximum):

@AmazingR

OpenStudy (anonymous):

34816+388 x=660 x find x

OpenStudy (mikeymaximum):

@JordanMayling

OpenStudy (anonymous):

can you solve above equation?

OpenStudy (anonymous):

660 x-388 x=34816 272 x=34816 x=34816/272=?

OpenStudy (mikeymaximum):

@Nerd94

OpenStudy (anonymous):

I don't have a calculator sooooo

OpenStudy (anonymous):

does anybody have the answer to this question